OpenStudy (anonymous):

Let f(x) = x2 - 3 and g(x) = 4x - 3, find (g.f)(-1) and (g.f)(2)

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

can u help me with another problem?

OpenStudy (anonymous):

Find (h.f.g)(-2) if f(x) = x2 - 5, g(x) = 3x + 1, h(x) = x + 3

OpenStudy (anonymous):

sorry bro will do but can u help me with a few more questions

OpenStudy (anonymous):

i did u see it?